# Generalized Hausdorff dimension of irrationals with Lagrange value exactly 3

### v1

@@ 1. Introduction
where $h:(0,\infty)\to(0,\infty)$ is an nondecreasing and right continuous function such that $h(r)\to 0$ as $r\to 0$. Such a function is called a dimension function or gauge function. See [11] for a comprehensive treatment of these (outer) measures.
@@
It is well known that $k^{-1}(3)$ is an uncountable dense set and it is a consequence of [7, Theorem 1] that $k^{-1}(3)$ has Hausdorff dimension 0. Indeed, for $t\in\mathbb{R}$ define
(1) $K_{t}=\{[0;a_{1},\dots,a_{n},\dots]\ \mid\ \text{there exists $(a_{-n})_{n\geq 0}\in(\mathbb{N}^{*})^{\mathbb{N}}$ such that}\\ [a_{k};a_{k+1},\dots,]+[0;a_{k-1},a_{k-2},\dots]\leq t,\forall k\in\mathbb{Z}\}.$
@@
A more general property of Liouville numbers is that they are immeasurable, that is, any translation invariant Borel measure on $\mathbb{R}$ either assigns zero measure to $\mathbb{L}$ or is not $\sigma$–finite on $\mathbb{L}$ [3, Theorem 1.1] (note that this includes the Hausdorff hh–measures). The main result of [3] applies to a broad class of sets, however, it does not apply to $k^{-1}(3)$ because of several reasons. Recall that set of Liouville numbers $\mathbb{L}$ forms a $G_{\delta}$ set that is invariant by rational translations (it is not difficult to see that $k^{-1}(3)$ is not invariant by such translations). A big difference is that from the topological point of view, the Liouville numbers are also larger than $k^{-1}(3)$.
@@
A forthcoming work of the first two authors with Zhe Cao will imply the following result: for any $x=[x_{0};x_{1},x_{2},\dots]\in k^{-1}(3)$ such that the inequality $\left|x-\frac{p}{q}\right|<\frac{1}{3q^{2}}$ has only finitely many solutions, there is an $N\in\mathbb{N}$ such that $0;x_{N},x_{N+1},\dots]\in K_{3}$. In particular, if we denote by $B_{3}$ the set of non–attainable elements of $k^{-1}(3)$, then for example for the dimension function $h(x)=\frac{1}{|\log\varepsilon|^{4}}$ one has that $\mathcal{H}^{h}(k^{-1}(3))=\infty$ but $\mathcal{H}^{h}(B_{3})=0$. In conclusion, there is a large gap between the generalized Hausdorff dimension of these sets, so the set of non–attainable elements $x\in k^{-1}(3)$ is also significantly smaller than $k^{-1}(3)$.
@@ 2. Preliminaries
Given $\underline{a}\in(\mathbb{N}^{*})^{\mathbb{Z}}$ define
@@
where $\gamma_{n+1}=[a_{n+1};a_{n+2},\dots]$ and $\eta_{n+1}=[0;a_{n},a_{n-1},\dots,a_{1}]$. In particular we have the formula $k(x)=\limsup_{i\to\infty}\lambda_{i}(x)$.
@@
Given a finite word of positive integers $w=(a_{1},\dots,a_{n})\in(\mathbb{N}^{*})^{n}$, define the closed sub-interval of $[0,1]$
@@
$\Sigma_{t}=\left\{\underline{a}\in(\mathbb{N}^{*})^{\mathbb{Z}}\ \mid\ \sup_{n\in\mathbb{Z}}\lambda_{n}(\underline{a})\leq t\right\}.$
@@
The subshift $\Sigma_{3}$ was characterized in [10]: after substituting $22\mapsto a$ and $11\mapsto b$ it coincides with the subshift of all Sturmian bi-infinite words (eventually periodic and aperiodic). It is well known that the closure of any aperiodic Sturmian bi-infinite word gives a minimal infinite subshift, so in particular, any orbit in that closure is an accumulation point. On the other hand, the eventually periodic Sturmian bi-infinite words can be approximated by the aperiodic (in the shift topology), thus any orbit of $\Sigma_{3}$ is an accumulation point. In particular the shift $\Sigma_{3}$ is homeomorphic to a Cantor space and consequently the set $K_{3}$ is a Cantor subset of $\mathbb{R}$.
@@
We define $\Sigma(t,n)$ to be the set of length-$n$ subwords of sequences in $\Sigma_{t}$. From [4, Theorem 1.1] we have the following:
@@
In fact, it is possible to prove that $|\Sigma(3,n)|\leq 9n^{3}$ for all $n\geq 1$, see [4, Corollary 3.13].
@@ 3. Dimensionality of $k^{-1}(3)$
Dimension functions $h$ for which $\mathcal{H}^{h}(h^{-1}(3))=\infty$ will be obtained by constructing suitable Cantor subsets $C\subset k^{-1}(3)$. In view of this, we introduce the following lemma concerning the fractal geometry of Cantor sets. The next lemma is a slightly modified version of [5, Lemma 1]. The proof is similar, and we include it here for completeness.
@@
$k^{-1}(3)\subset\bigcup_{n_{0}\geq 2}\bigcup_{(a_{1},\dots,a_{n_{0}-1})\in(\mathbb{N}^{*})^{n_{0}-1}}\{[0;a_{1},\dots,a_{n_{0}-1},\gamma]:\gamma\in K_{3+6^{-m}}\}.$
@@
It is proved in [9, Lemma 2.2] that for all dimension functions $h$ and all sets $E\subset\mathbb{R}$, we have:
@@
4. The set $K_{3}$

### v2

@@ 1. Introduction
where $h:(0,\infty)\to(0,\infty)$ is a nondecreasing and right continuous function such that $h(r)\to 0$ as $r\to 0$. Such a function is called a dimension function or gauge function. See [16] for a comprehensive treatment of these (outer) measures.
@@
It is well known that $k^{-1}(3)$ is an uncountable dense set and it is a consequence of [10, Theorem 1] that $k^{-1}(3)$ has Hausdorff dimension 0. Indeed, for $t\in\mathbb{R}$ define
(1) $K_{t}=\{[0;a_{1},\dots,a_{n},\dots]\ \mid\ \text{there exists $(a_{-n})_{n\geq 0}\in(\mathbb{N}_{>0})^{\mathbb{N}}$ such that}\\ [a_{k};a_{k+1},\dots,]+[0;a_{k-1},a_{k-2},\dots]\leq t,\forall k\in\mathbb{Z}\}.$
@@
A more general property of Liouville numbers is that they are immeasurable, that is, any translation invariant Borel measure on $\mathbb{R}$ either assigns zero measure to $\mathbb{L}$ or is not $\sigma$–finite on $\mathbb{L}$ [4, Theorem 1.1] (note that this includes the Hausdorff hh–measures). The main result of [4] applies to a broad class of sets, however, it does not apply to $k^{-1}(3)$ because of several reasons. Recall that set of Liouville numbers $\mathbb{L}$ forms a $G_{\delta}$ set that is invariant by rational translations (it is not difficult to see that $k^{-1}(3)$ is not invariant by such translations). A big difference is that from the topological point of view, the Liouville numbers are also larger than $k^{-1}(3)$.
@@
    $\limsup_{\varepsilon\to 0}\frac{|\log h(\varepsilon)|}{\log|\log\varepsilon|}<3,$
    then $\mathcal{H}^{h}(K_{3})=\infty$.
  2. (2)
    If
    $\limsup_{\varepsilon\to 0}\frac{|\log h(\varepsilon)|}{\log|\log\varepsilon|}>3,$
@@
In fact the previous theorem is corollary of a stronger result.
Theorem 1.5.
For the dimension function
$g(\varepsilon)=\frac{\log|\log\varepsilon|}{|\log\varepsilon|^{3}(\log\log|\log\varepsilon|)},$
we have that $\mathcal{H}^{g}(K_{3})<\infty$.
For a large enough constant $\hat{C}>0$, the dimension function
$g(\varepsilon)=\frac{\exp(\hat{C}\sqrt{\log|\log\varepsilon|})}{|\log\varepsilon|^{3}}$
satisfies $\mathcal{H}^{g}(K_{3})=\infty$.
Then for example for the dimension function $h(\varepsilon)=\frac{1}{|\log\varepsilon|^{3}}$ one has that $\mathcal{H}^{h}(k^{-1}(3))=\infty$ but $\mathcal{H}^{h}(K_{3})=0$. In conclusion, there is a large gap between the generalized Hausdorff dimension of these sets, so the set $K_{3}$ is significantly smaller than $k^{-1}(3)$.
Denote by $B_{3}$ the set of non–attainable elements of $k^{-1}(3)$, that is
$B_{3}=\left\{x\in\mathbb{R}\setminus\mathbb{Q}:\left|x-\frac{p}{q}\right|<\frac{1}{3q^{2}}\text{ has only finitely many solutions }\frac{p}{q}\in\mathbb{Q}\right\}.$
A recent work [3] by the first two authors with Zhe Cao imply the following result: for any $x=[x_{0};x_{1},x_{2},\dots]\in B_{3}$, there is an $N\in\mathbb{N}$ such that $0;x_{N+1},x_{N+2},\dots]\in K_{3}$. However, while the Cantor set $K_{3}$ is forward invariant by the Gauss map, the set $B_{3}$ is not. The next theorem shows that the non–attainable elements $B_{3}$ are even smaller than $K_{3}$ and thus smaller than $k^{-1}(3)$.
Theorem 1.6.
Let $h$ be a dimension function.
  1. (1)
    If
    $\limsup_{\varepsilon\to 0}\frac{|\log h(\varepsilon)|}{\log|\log\varepsilon|}<2,$
    then $\mathcal{H}^{h}(B_{3})=\infty$.
  2. (2)
    If
    $\limsup_{\varepsilon\to 0}\frac{|\log h(\varepsilon)|}{\log|\log\varepsilon|}>2,$
    then $\mathcal{H}^{h}(B_{3})=0$.
The proof of the previous theorem is essentially the same as Theorem 1.4. The key distinction is that, in the case of $K_{3}$, one must account for permutations of renormalizable words, reflecting its invariance under the Gauss map. By contrast, the structure of $B_{3}$ is considerably more rigid so this additional counting is not required. For completeness, one could also formulate an analogue of Theorem 1.5 for $B_{3}$, but we refrain from doing so here in order to avoid unnecessary repetition. From this version it follows that $\mathcal{H}^{h}(B_{3})=0$ for the gauge function $h(\varepsilon)=\frac{1}{|\log\varepsilon|^{2}}$.
@@ 2.1. Continued fraction and Hausdorff dimension
Given $\underline{a}\in(\mathbb{N}_{>0})^{\mathbb{Z}}$ define
@@
where $\gamma_{n+1}=[a_{n+1};a_{n+2},\dots]$ and $\eta_{n+1}=[0;a_{n},a_{n-1},\dots,a_{1}]$. In particular we have the formula $k(x)=\limsup_{n\to\infty}\lambda_{n}(x)$.
@@
Given a finite word of positive integers $w=(a_{1},\dots,a_{n})\in(\mathbb{N}_{>0})^{n}$, define the closed sub-interval of $[0,1]$
@@
The well known estimate of Jarnik [7] says that the Hausdorff dimension of the Gauss-Cantor set $C(n)=\{[0;a_{1},a_{2},\dots]:1\leq a_{i}\leq n,\forall i\geq 1\}$ satisfies $1-C_{2}/n<\dim_{H}(C(n))<1-C_{1}/n$ for some explicit constants $0<C_{1}<C_{2}$ and all $n$ large. For our purposes, Jarnik result is enough, however a much more precise asymptotic estimate is due to Hensley [6]:
(5) $\dim_{H}(C(n))=1-\frac{6}{\pi^{2}n}-\frac{72\log n}{\pi^{4}n^{2}}+O\left(\frac{1}{n^{2}}\right).$
Another related result by Hensley estimates the cardinality of the set
[…]
Hensley [6] proved that for any $n>1$, there is a constant $E_{n}>0$ such that $\lim_{x\to\infty}x^{-2\dim_{H}(C(n))}|Q_{n}(x)|=E_{n}$. For our purposes, we will need a more elementary estimate that is not dependent on the rate of convergence.
Proposition 2.1.
For all $x>1$,
(6) $x^{2\dim_{H}(C(n))}\leq Q_{n}(x)\leq 2x^{2\dim_{H}(C(n))}.$
Proof.
Recall that a Gauss-Cantor set $K(B)$ is defined by an alphabet $B=\{\beta_{1},\dots,\beta_{\ell}\}$ and for each word $\beta_{j}\in(\mathbb{N}_{>0})^{r_{j}}$, the interval $I_{j}=I(\beta_{j})$ and $\psi|_{I_{j}}:=G^{r_{j}}|_{I_{j}}$ is an iterate of the Gauss map $G(x)=\frac{1}{x}-\lfloor\frac{1}{x}\rfloor$. This defines an expanding map $\psi\colon I_{1}\cup\dotsb\cup I_{\ell}\to I$ with Markov partition $I_{1},\dots,I_{\ell}$. According to Palis–Takens [13, Chapter 4, Pages 68-71], if we let
$\lambda_{j}=\inf|\psi^{\prime}|_{I_{j}}|,\qquad\Lambda_{j}=\sup|\psi^{\prime}|_{I_{j}}|$
and $d_{1},d_{2}\geq 0$ be such that
$\sum_{i=1}^{\ell}\lambda_{j}^{-d_{2}}=1,\qquad\sum_{i=1}^{\ell}\Lambda_{j}^{-d_{1}}=1,$
then
(7) $d_{1}\leq\dim_{H}(K(B))\leq d_{2}.$
Let us discuss how to find estimates for $d_{1}$ and $d_{2}$. The iterates of the Gauss map are given explicitly by
$\psi|_{I_{j}}(x)=\dfrac{q^{(j)}_{r_{j}}x-p^{(j)}_{r_{j}}}{p^{(j)}_{r_{j}-1}-q^{(j)}_{r_{j}-1}x}$
where $\dfrac{p^{(j)}_{k}}{q^{(j)}_{k}}=[0;b^{(j)}_{1},\dots,b^{(j)}_{k}$] and $\beta_{j}=(b^{(j)}_{1},\dots,b^{(j)}_{r_{j}})$.
Hence
$(\psi|_{I_{j}})^{\prime}(x)=\frac{(-1)^{r_{j}-1}}{(p^{(j)}_{r_{j}-1}-q^{(j)}_{r_{j}-1}x)^{2}}.$
Let $x=[c_{0},c_{1},c_{2},\dots]$ and $\frac{p_{n}}{q_{n}}=[c_{0},c_{1},\dots,c_{n}]$. Then
$\frac{1}{2q_{n}q_{n+1}}<\frac{1}{q_{n}(q_{n}+q_{n+1})}<\left|x-\frac{p_{n}}{q_{n}}\right|<\frac{1}{q_{n}q_{n+1}},$
and therefore
$\frac{1}{2q_{n+1}}<|q_{n}x-p_{n}|<\frac{1}{q_{n+1}}.$
Therefore,
$(q_{r_{j}}^{(j)})^{2}<|(\psi|_{I_{j}})^{\prime}(x)|=\frac{1}{(p^{(j)}_{r_{j}-1}-q^{(j)}_{r_{j}-1}x)^{2}}<(2q_{r_{j}}^{(j)})^{2}.$
Thus
$(q_{r_{j}}^{(j)})^{2}\leq\lambda_{j}=\inf|\psi^{\prime}|_{I_{j}}|\leq\Lambda_{j}=\sup|\psi^{\prime}|_{I_{j}}|\leq(2q_{r_{j}}^{(j)})^{2}.$
Consider the set $P_{n}(x,m)=\{(a,b)\in Q_{n}(x):x/2^{m/2}<b\leq x/2^{(m-1)/2}\}$. Notice that writing $a/b=[0;b_{1},\dots,b_{\ell}]$, each $0;b_{1},\dots,b_{i}]$ belongs to a different $P_{n}(x,m)$ for $1\leq b_{i}\leq\ell$. In particular the intervals $I(b_{1},\dots,b_{\ell})$ where $(a,b)\in P_{n}(x,m)$ form a Markov partition of the Gauss-Cantor set $C(n)$. By (7) we have that
$1=\sum_{i=1}^{|P_{n}(x,m)|}\lambda_{j}^{-d_{2}}\leq|P_{n}(x,m)|x^{-2d_{2}}2^{d_{1}m}\leq|P_{n}(x,m)|x^{-2\dim_{H}(C(n))}2^{m},$
whence
$|Q_{n}(x)|=\sum_{m\geq 1}|P_{n}(x,m)|\geq x^{2\dim_{H}(C(n))}.$
Similarly
$1=\sum_{i=1}^{|P_{n}(x,m)|}\Lambda_{j}^{-d_{1}}\geq|P_{n}(x,m)|x^{-2d_{1}}2^{d_{2}(m+1)}\geq|P_{n}(x,m)|x^{-2\dim_{H}(C(n))}2^{m+1},$
whence whence
$|Q_{n}(x)|=\sum_{m\geq 1}|P_{n}(x,m)|\leq 2x^{2\dim_{H}(C(n))}.$
2.2. The subshift $\Sigma_{3}$
@@
$\Sigma_{t}=\left\{\underline{a}\in(\mathbb{N}_{>0})^{\mathbb{Z}}\ \mid\ \sup_{n\in\mathbb{Z}}\lambda_{n}(\underline{a})\leq t\right\}.$
@@
The subshift $\Sigma_{3}$ was characterized in [14]: after substituting $22\mapsto a$ and $11\mapsto b$ it coincides with the subshift of all balanced bi-infinite words (eventually periodic and aperiodic). It is well known that the closure of any aperiodic Sturmian bi-infinite word gives a minimal infinite subshift, so in particular, any orbit in that closure is an accumulation point. On the other hand, the eventually periodic balanced bi-infinite words can be approximated by the aperiodic (in the shift topology), thus any orbit of $\Sigma_{3}$ is an accumulation point. In particular the shift $\Sigma_{3}$ is homeomorphic to a Cantor space and consequently the set $K_{3}$ is a Cantor subset of $\mathbb{R}$.
The shift $\Sigma_{3}$ can be described via the SS–adic representation of balanced sequences. Background material and proofs of the statements below can be found in [15]. We call the substitutions $U$ and $V$ the inner renormalization operators, defined by
[…]
Let $(\alpha,\beta)$ be a pair of finite words over the alphabet $\{a,b\}$. Define the exterior renormalization operators $\overline{U},\overline{V}$ by
$\overline{U}:(\alpha,\beta)\mapsto(\alpha\beta,\beta),\quad\overline{V}:(\alpha,\beta)\mapsto(\alpha,\alpha\beta).$
The Cohn’s tree is a binary tree of finite words over $\{a,b\}$ defined as follows: take as root the word $ab$ and given a vertex, its children in the next layer are obtained by applying $U$ and $V$. Let $P$ be the set of vertices of Cohn’s tree. These finite words together with $\{a,b\}$ are called (lower) Christoffel words. It is a fact that a subword is subfactor of a balanced sequence if and only if is subfactor of some Christoffel word if and only if is a subfactor of some Sturmian sequence. By [5, Lemma 3.11] we have a control of how large is this Christoffel word.
Lemma 2.2.
Let $w$ be a factor of a word in $P$. Then the length of the shortest word in $P$ containing $w$ is strictly smaller than $3|w|$.
Analogously, one can define a tree of alphabets from word pairs as follows: the root is $(a,b)$ and given a vertex, its children in the next layer are obtained by applying $\overline{U}$ and $\overline{V}$. The set of vertices of this tree $\overline{P}$ are also called Christoffel pairs.
The Cohn’s tree and the tree of alphabets are related through the concatenation operation: given an alphabet $(\alpha,\beta)\in\overline{P}$ we have that $\alpha\beta\in P$ and conversely, given any Christoffel word $w\in P$, there is a unique factorization $w=\alpha\beta$ where $(\alpha,\beta)\in\overline{P}$. Moreover, given any $(\alpha,\beta)\in\overline{P}$, there are some substitutions $R_{1},\dots,R_{n}$ such that $\alpha=(R_{1}\dots R_{n})(a)$ and $\beta=(R_{1}\dots R_{n})(b)$.
Another useful property of lower Christoffel words $P$ is that they are completely determined by counting their letters. More precisely, for $w\in P$, the Frobenius coordinates $(|w|_{a}+|w|_{b},|w|_{b})$ uniquely determine $w$, where $|w|_{c}$ denotes the number of occurrences of the letter $c\in\{a,b\}$. Furthermore, every word in $P$ is primitive, that is, no $w\in P$ is a proper power of a shorter word. Consequently, all cyclic permutations of a given $w$ are distinct. In addition, cyclic permutations of two distinct words $w,w^{\prime}\in P$ cannot coincide, since their Frobenius coordinates are different.
Continued fractions are closely related to the Christoffel pairs. Indeed, for each continued fraction $0;a_{1},\dots,a_{n}]$ we can associate a unique Christoffel pair as follows. Define $(\alpha_{0},\beta_{0})=(a,b)$, $(\alpha_{1},\beta_{1})=\overline{V}^{a_{1}-1}(\alpha_{0},\beta_{0})$, $(\alpha_{2},\beta_{2})=\overline{U}^{a_{2}}(\alpha_{1},\beta_{1})$, $(\alpha_{3},\beta_{3})=\overline{V}^{a_{3}}(\alpha_{2},\beta_{2})$ and so on. In particular we will have that $q_{0}=|\alpha_{0}|/2$, $q_{1}=|\beta_{1}|/2$, $q_{2}=|\alpha_{2}|/2$, $q_{3}=|\beta_{3}|/3$ and so on, where $q_{1},\dots,q_{n}$ are the denominators of the convergents of $0;a_{1},\dots,a_{n}]$. This process can be imagined as the mediant process used to build the convergents of the continued fraction $0;a_{1},\dots,a_{n}]$.
$\alpha$$\beta$$\alpha\beta$$\alpha\beta^{2}$$\alpha\beta^{3}$$\alpha\beta^{4}\dots$
$\alpha$$\beta$$\alpha\beta^{3}\dots$$\alpha\beta^{3}((\alpha\beta^{3})^{2}\beta)^{4}$$(\alpha\beta^{3})^{2}\beta$
Figure 1. Example of how constructing alphabets is similar to build the convergents of a number $[0;1,3,2,4]$.
Recall that given fractions $0<\frac{p}{q}<\frac{r}{s}$ the mediant is
$\frac{p}{q}\oplus\frac{r}{s}=\frac{p+r}{q+s}.$
Given $n\geq 1$, the Farey sequence $\mathcal{F}_{n}$ of order $n$ is the set of reduced coprime fractions $0\leq\frac{p}{q}\leq 1$ with $1\leq q\leq n$ increasingly ordered. Two fractions $\frac{p}{q}<\frac{r}{s}$ are Farey neighbors in some $\mathcal{F}_{n}$ if and only if psqr=1ps-qr=1 and $\max\{q,s\}\leq n<q+s$. The mediant $\frac{p+r}{q+s}$ will appear between $\frac{p}{q}$ and $\frac{r}{s}$ precisely at term $\mathcal{F}_{q+s}$.
The rule used to construct the tree of alphabets is related to the mediant rule used in the construction of the Farey sequence. We can build a bijection $\Theta$ from the Christoffel words $P\cup\{a,b\}$ to the positive fractions $\mathbb{Q}\cap[0,1]$ as follows. Given $w\in P\cup\{a,b\}$, we define $\Theta(w)=\frac{|w|_{b}}{|w|_{a,b}}$ where $|w|_{a,b}$ is the length of $w$ over the alphabet $\{a,b\}$ and $|w|_{b}$ is the number of letters $b$ in $w$. The following fact is consequence of [17, Lemma 3.1].
Lemma 2.3.
For each Farey neighbors $\frac{p}{q}<\frac{r}{s}$, there is a unique pair of words $(\alpha,\beta)\in\overline{P}$ such that $\Theta(\alpha)=p/q$ and $\Theta(\beta)=r/s$.
@@
We define $\Sigma(t,n)$ to be the set of length-$n$ subwords of sequences in $\Sigma_{t}$. From [5, Theorem 1.1] we have the following:
@@
In fact, it is possible to prove that $|\Sigma(3,n)|\leq 9n^{3}$ for all $n\geq 1$, see [5, Corollary 3.13].
@@ 3. Dimensionality of $k^{-1}(3)$
Dimension functions $h$ for which $\mathcal{H}^{h}(h^{-1}(3))=\infty$ will be obtained by constructing suitable Cantor subsets $C\subset k^{-1}(3)$. In view of this, we introduce the following lemma concerning the fractal geometry of Cantor sets. The next lemma is a slightly modified version of [8, Lemma 1]. The proof is similar, and we include it here for completeness.
@@
$k^{-1}(3)\subset\bigcup_{n_{0}\geq 2}\bigcup_{(a_{1},\dots,a_{n_{0}-1})\in(\mathbb{N}_{>0})^{n_{0}-1}}\{[0;a_{1},\dots,a_{n_{0}-1},\gamma]:\gamma\in K_{3+6^{-m}}\}.$
@@
It is proved in [12, Lemma 2.2] that for all dimension functions $h$ and all sets $E\subset\mathbb{R}$, we have:
@@
4. Dimensionality of $K_{3}$
For the Cantor set
$K_{3}=\{[0;a_{1},\dots,a_{n},\dots]\ \mid\ \text{there exists $(a_{-n})_{n\geq 0}\in(\mathbb{N}_{>0})^{\mathbb{N}}$ such that}\\ [a_{k};a_{k+1},\dots,]+[0;a_{k-1},a_{k-2},\dots]\leq 3,\forall k\in\mathbb{Z}\},$
we obtain a reasonably explicit description of its generalized Hausdorff measures. However, due to additional combinatorial complications, the result is not as precise as in the case of $k^{-1}(3)$.
Recall that for any fixed $n\geq 1$, the set $\{I(w):w=(a_{1},\dots,a_{n})\in\Sigma(3,n)\}$ is a covering of $K_{3}$ by closed intervals.